# Language-Based Bayesian Optimization Research Assistant (BORA)

### v1

@@ 3.4.1 Action Selection
BORA’s policy helps it to make informed choices about engaging the LLM without relying on data-hungry RL algorithms, thus maintaining BORA’s practicality and effectiveness in real-life scenarios. The optimization starts with the Vanilla BO action $a_{1}$, and the subsequent action selection depends on (a) the average uncertainty $\sigma_{\text{mean}}^{\text{GP}}$ of the common GP over the search space $\mathcal{X}$ to determine the necessity and type of LLM intervention, and (b) the BO performance detection, as well as the performance success (or trust in) of the previous LLM interventions. When the GP’s uncertainty is high and above a pre-defined threshold ($\sigma_{t,\text{mean}}^{\text{GP}}>\sigma_{t,\text{upper}}$), BO needs significant guidance from the LLM, triggering a complete ‘take-over’ by the LLM in the search, suggesting new points informed by its own internal reasoning mechanism. As the GP’s uncertainty decreases ($\sigma_{t,\text{lower}}\leq\sigma_{t,\text{mean}}^{\text{GP}}\leq\sigma_{t,\text{upper}}$), the LLM becomes less involved by relying only on BO suggested points, but still using its ICL capacity based on both $\mathcal{D}$ and $\mathcal{C}$ to select the most promising ones. When the GP’s uncertainty is low enough ($\sigma_{t,\text{mean}}^{\text{GP}}<\sigma_{t,\text{lower}}$), BO has a better approximation of objective function’s landscape and no longer needs guidance from the LLM. The rationale behind remark (b) is that the LLM should gain more trust as the LLM suggestions exhibit better performance, which in turn triggers the plateau duration to be re-defined as shorter, allowing the LLM to intervene more frequently. Conversely, if the LLM’s so-far observed performance declines, its trust in itself diminishes and, consequently, its interventions are reduced, which results in longer plateau duration adjustments before invoking its assistance. In short, the action selection at each step $t$ follows the policy $\pi$ described below, where the GP parameters are updated after every action accordingly:
@@
  • Else if $\sigma_{t,\text{mean}}^{\text{GP}}>\sigma_{t,\text{upper}}$to action $a_{2}$,
@@ 3.4.2 Selection Mechanism
The above action selection is realized by calculating and updating in every step the uncertainties from a set of fixed $q$ monitoring points $x_{\text{mon}}^{(i)}$ that were randomly sampled before the optimization starts. Specifically,
@@
where $\sigma_{t}(\cdot)$ represents the uncertainty of the GP at a given point in iteration $t$.
@@
where $y_{t}^{\text{max}}=\max\bigl{(}\{y|(x,y)\in\mathcal{D}_{t}\}\bigr{)}$. That is, if for the past $m$ consecutive BO steps, there is not enough performance improvement (w.r.t a set percentage $\gamma)$, then the LLM involvement is triggered. The plateau duration $m$ is initialized at $m_{\text{init}}=\lceil 2\sqrt{d}\rceil$, empirically set to vary between $m_{\text{min}}=0$ and $m_{\text{max}}=3m_{\text{init}}$, and is automatically adjusted at every LLM intervention step $l$ (here $l$ counts the number of times actions $a_{2}$ or $a_{3}$ are invoked). The adjustment depends on the current ‘trust’ $T_{l}\in[0,1]$ BORA has on the LLM, which in turn relies on the LLM performance observed so far. Specifically
@@
$\text{score}_{\text{interv}}^{(l)}$ =$=\begin{cases}1,&\text{if }r_{l}>0,\\ \frac{1}{1+e^{-\frac{r_{l}}{|y_{t-1}^{\text{max}}|+\epsilon}}},&\text{if }r_{l}\leq 0,\end{cases}$ (9)
@@
3:  Initialize policy parameters $\sigma_{0,\text{mean}}^{\text{GP}}$, $\sigma_{0,\text{max}}^{\text{GP}}$, $\sigma_{0,\text{upper}}$, $\sigma_{0,\text{lower}}$, $m$, =$\mathcal{H}=\{0.9\}$, $\gamma=0.05$;
@@
11:        $a=a_{3}$$\mathcal{S}_{t}=\{x_{t}^{(k)}\}_{k=1}^{\min(2,i_{\text{max}}+n_{\text{init}}-i)}$;
@@ 4 Experiments
We validated BORA’s performance against current state-of-the-art methods on both synthetic functions and various real-world tasks, with dimensionality ranging from 4 to 15 independent variables. Section 4.1 outlines the experimental setup while Section 4.2 highlights the results. Details on the benchmarks, the method implementations, and the reproducibility details can be found in the SM. The source code is available at https://anonymous.4open.science/r/bora-the-explorer.
@@ 4.1.1 Synthetic Function Benchmarks
  • Levy (10D): A function with a highly rugged landscape. All inputs are bounded in $[-10,10]$ with the maximum at $[1,\dots,1]$.
  • Ackley (15D): A challenging high dimensional function with several local maxima. Input bounds are $[-30,20]$ with the maximum at $[0,\dots,0]$.
@@ 4.1.3 BORA
We implemented BORA using OpenAI’s most cost-effective model, GPT-4o-mini, which was not fine-tuned in our effort to make BORA more accessible to users with limited resources. For the BO action implementation, the GP uses a Matérn kernel, and the acquisition function is Expected Improvement.
@@ 4.1.4 Baselines
For ColaBO and HypBO, to avoid the impracticality of relying on humans to provide inputs for multiple trials across all experiments, we used the LLM GPT-4o-mini to generate the ‘human’ inputs. Likewise, we used the same task description prompts as used for BORA, to ensure consistency. For HypBO on the Hydrogen Production experiment, we employed the most realistic hypothesis used in Cissé et al. (2024), namely ‘What They Knew’, which encapsulates any human knowledge available prior to the execution of the experiment, thus precluding any post-experiment hindsight.
@@ 4.1.5 Experimental Protocol
The optimization performance was measured using the maximum objective value found so far, the cumulative regret, and statistical tests to measure significance. The maximum number of samples was set to 105 to account for realistic budgets with expensive functions. Average results of 10 repeated trials with distinct random seeds are reported. All methods were initialized with 5 initial samples apart from LAEA, for which we used 15 initial samples to keep the same number of evaluations to population size ratio as in Hao et al. (2024).
Figure 5: BORA vs Baselines on six experiments. Solid lines show average values while shaded areas indicate ±$\pm 0.25$ standard error. For visual clarity, some plots are zoomed in to show results up to 60 iterations, as the trends mostly stabilize beyond this point. Full results in SM.
@@ 4.2.1 Synthetic Functions
Figure 3 illustrates the exploration strategy of BORA on the Branin (2D) function. The LLM interventions helped to uncover two out of the three possible locations of the global maximum of Branin. This is further illustrated in Figure 5, which shows that BORA significantly outperforms the baseline comparisons for the higher-dimension functions Levy (10D) and Ackley (15D). A key advantage of BORA is its LLM-informed initial sampling. For mathematical functions, BORA systematically suggests initializing at critical points such as the edges, central points, or other remarkable points like $[0,\dots,0]$. This strategy is particularly well-suited for the Levy function, whose search space bounds are symmetric, and its optimum is at $[1,\dots,1]$, almost always converging in its initialization stage. However, that strategy is less beneficial on the Ackley function because its bounds are asymmetric. Despite that, the LLM’s ability to reflect and learn from the previous samples appears to help mitigate any unfavorable initializations. While HypBO demonstrates a similar benefit through its initial sampling in hypothesized regions, its performance is comparatively weaker because it relies on random initial sampling within these regions, resulting in less effective exploration of the search space. For ColaBO, which only works with a single input prior, the prior tended to be around one of the edges, which overall limits its convergence speed. Additionally, the left panel of Figure 4 shows how BORA’s iterative hypothesis generation, informed by previous data, helps mitigate stagnant optimization, and discards the LLM when it is no longer needed. Notably, a sharp drop in the GP uncertainty is evident when vanilla BO is used due to is proved exploration-exploitation guarantees, as opposed to the less rigorous LLM where the uncertainty is bound to its inherent sampling strategy. Nevertheless, the dynamic synergistic effect of BO coupled with updated LLM hypotheses allows for faster convergence overall, in comparison to other baselines as illustrated in the last two bar plots of Figure 6.
@@ 4.2.2 Real-World Applications
BORA also exhibits superior performance across diverse real-world optimization problems, following the trends observed in the synthetic benchmarks. As shown in Figure 5, while BORA’s initial sampling is on par with other input-based baselines for the Solar Energy and Sugar Beet Production experiments, its overall performance on all experiments significantly surpasses the baselines. This is particularly evident in the 7D Pétanque experiment, where BORA’s diverse initial hypotheses based on trajectory dynamics led to a remarkable gain in score of 35 in the early stages compared to the baselines. This knowledge and context-based input bridges the knowledge gap typically encountered in early-stage optimization, providing BORA with a critical advantage, as shown in Figure 6. A similar effect is also observed in the Hydrogen Production experiment as illustrated in the right panel of Figure 4. In addition, in the later stage of the optimization the LLM further pushes the optimization to uncover new optima after the progress had stalled, thus gaining more trust. The increasing trend in the GP uncertainty here is a side-effect of the continual interventions of the LLM which translates to a rather explorative and less exploitative strategy based on its inherent domain reasoning around cumulatively accrued scientific data. This goes beyond some of the near-instant convergence noted in the optimization of the synthetic functions because they are proximal to a special value. Figure 2 illustrates this by showing how the LLM reflects on the progress and generates hypotheses on the Sugar Beet Production experiment. While other baselines, particularly knowledge-based methods such as HypBO and local BO approaches such as TuRBO, demonstrate improved performance as the optimization progresses and more data is gathered, they often struggle to match BORA’s sustained performance as the 105-sample budget mark is approached. In the Hydrogen Production experiment, this adaptive strategy ultimately achieved a 47% reduction in cumulative regret compared to ColaBO, demonstrating BORA’s faster convergence and robustness in navigating complex, high-dimensional search spaces. To assess the significance of the performance difference w.r.t mean cumulative regret between BORA and its best two competitors, we performed a sign test which revealed that BORA performs consistently better than HypBO with a p-value of 0.02 at a 95% confidence level with a Bonferroni correction Bonferroni (1936), but not against ColaBO with a p-value of 0.20, yet still outscoring it in 5 out of 6 tasks. The superior performance of the hybrid approach in BORA was further validated by ablation studies that used only the LLM (action $a_{2}$) for optimizing Hydrogen Production (10D) and the Ackley function (15D) (see SM). While performing quite well in the initial stages for these two problems, the use of the LLM alone was ultimately less effective than the dynamic hybrid BO/LLM approach in BORA. It is important to emphasize that these results do not mean LLMs are ‘smarter’ than domain experts. Rather, they highlight BORA’s ability to update and refine its hypotheses based on new data, which is not possible in the HypBO implementation Cissé et al. (2024), while also fostering user engagement by generating real-time optimization progress commentary and a final summary report (see SM). One potential limitation of BORA, however, is the stochastic nature of the LLM reasoning, which can diverge considerably even with identical prompts.
@@ A.0.1 Main results
Figure 7: Full results of the BORA vs Baselines on the six experiments. Solid lines show average values while shaded areas indicate ±$\pm 0.25$ standard error.
@@
To further evaluate BORA’s contributions, we compared it with an LLM-only optimizer, which effectively isolates the LLM component of BORA and functions as an independent optimizer. This comparative analysis allowed us to determine the degree to which the incorporation of Bayesian Optimization (BO) principles enhances performance beyond the capabilities of the LLM alone. LLM-Only uses identical prompts, albeit with minor modifications in the role-based prompting mechanism to ensure that it acknowledges its status as the sole optimizer without anticipating BO interventions. Furthermore, it employs the same hypothesis generation mechanism as BORA’s LLM action $a_{2}$, relying exclusively on the outputs of the LLM for optimization. Additionally, LLM-Only uses the same reflection and fallback mechanisms. In the rare event where the LLM fails to produce a valid hypothesis, the corresponding hypothesis point is replaced with a randomly generated point to maintain continuity.
@@
We evaluated BORA and LLM-Only on the more complex synthetic function Ackley (15D) and on the most demanding highest-dimensional real-world task, Hydrogen Production (10D), and reported the maximum objective value found. To ensure adherence to the experimental protocol outlined in the manuscript, LLM-Only was executed with an optimization budget of 105 samples and repeated 10 times.
@@ Results
Figure 8 shows the results of this comparison. LLM-Only shows competitive performance during the early stages of optimizing both Ackley (15D) and Hydrogen Production (10D). This achievement can be attributed to LLMs’ ability to perform inductive biases, which are used during hypothesis generation. However, LLM-Only fails to sustain its initial performance and quickly stagnates, falling behind BORA significantly in the later stages of the optimization. This stagnation highlights the limitations of relying solely on LLM inference to optimize high-dimensional spaces. In particular, LLMs lack the balanced exploration-exploitation trade-offs that are inherent to BO-based methods, which are achieved through posterior uncertainty modeling. In contrast, BORA, as a hybrid BO/LLM approach, dynamically merges structured sampling from BO with the LLM’s contextual understanding, facilitating sustained improvements as the optimization progresses. Notably, BORA achieved up to a 67% improvement in Hydrogen Production and a 31% improvement in the Ackley benchmark based on the maximum objective values discovered. While BORA considerably outperforms LLM-Only, it is interesting to note that while LLM-Only struggles to sustain its initial, early-stage performance, it still outperforms half of the baselines studied here in this small-budget setting.

### v2

@@ 3.4.1 Action Selection
BORA’s policy helps it to make informed choices about engaging the LLM without relying on data-hungry RL algorithms, thus maintaining BORA’s practicality and effectiveness in real-life scenarios. The optimization starts with the Vanilla BO action $a_{1}$, and the subsequent action selection depends on (a) the average uncertainty $\sigma_{\text{mean}}^{\text{GP}}$ of the common GP over the search space $\mathcal{X}$ to determine the necessity and type of LLM intervention, and (b) the BO performance plateau detection, as well as the performance success (or trust in) of the previous LLM interventions. When the GP’s uncertainty is high and above a pre-defined threshold ($\sigma_{t,\text{mean}}^{\text{GP}}>\sigma_{t,\text{upper}}$), BO needs significant guidance from the LLM, triggering a complete ‘take-over’ by the LLM in the search, suggesting new points informed by its own internal reasoning mechanism. As the GP’s uncertainty decreases ($\sigma_{t,\text{lower}}\leq\sigma_{t,\text{mean}}^{\text{GP}}\leq\sigma_{t,\text{upper}}$), the LLM becomes less involved by relying only on BO suggested points, but still using its ICL capacity based on both $\mathcal{D}$ and $\mathcal{C}$ to select the most promising ones. When the GP’s uncertainty is low enough ($\sigma_{t,\text{mean}}^{\text{GP}}<\sigma_{t,\text{lower}}$), BO has a better approximation of objective function’s landscape and no longer needs guidance from the LLM. The rationale behind remark (b) is that the LLM should gain more trust as the LLM suggestions exhibit better performance, which in turn triggers the plateau duration to be re-defined as shorter, allowing the LLM to intervene more frequently. Conversely, if the LLM’s so-far observed performance declines, its trust in itself diminishes and, consequently, its interventions are reduced, which results in longer plateau duration adjustments before invoking its assistance. In short, the action selection at each step $t$ follows the policy $\pi$ described below, where the GP parameters are updated after every action accordingly:
@@
  • Else if $\sigma_{t,\text{mean}}^{\text{GP}}>\sigma_{t,\text{upper}}$ $\to$ action $a_{2}$,
@@ 3.4.2 Selection Mechanism
The above action selection is realized by calculating and updating in every step the uncertainties from a set of fixed $q$ monitoring points $x_{\text{mon}}^{(i)}$ that are randomly sampled before the optimization starts. Specifically,
@@
where $\sigma_{t}(\cdot)$ represents the uncertainty of the GP at a given point in iteration $t$. Here, the 50% and 30% fractions serve as empirically tuned bounds that consistently balance BO exploitation with LLM exploration across diverse tasks.
@@
where $y_{t}^{\text{max}}=\max\bigl{(}\{y|(x,y)\in\mathcal{D}_{t}\}\bigr{)}$. That is, if for the past $m$ consecutive BO steps, there is not enough performance improvement (w.r.t a set percentage $\gamma)$, then the LLM involvement is triggered. The plateau duration $m$ is initialized at $m_{\text{init}}=\lceil 2\sqrt{d}\rceil$, set to vary between $m_{\text{min}}=0$ and $m_{\text{max}}=3m_{\text{init}}$, and is automatically adjusted at every LLM intervention step $l$ (here $l$ counts the number of times actions $a_{2}$ or $a_{3}$ are invoked). The adjustment depends on the current ‘trust’ $T_{l}\in[0,1]$ BORA has on the LLM, which in turn relies on the LLM performance observed so far. Specifically
@@
$\text{score}_{\text{interv}}^{(l)}$ $=\begin{cases}1,&\text{if }r_{l}>0,\\ \frac{1}{1+e^{-\frac{r_{l}}{|y_{t-1}^{\text{max}}|+\epsilon}}},&\text{if }r_{l}\leq 0,\end{cases}$ (9)
@@
3: Initialize policy parameters $\sigma_{0,\text{mean}}^{\text{GP}}$, $\sigma_{0,\text{max}}^{\text{GP}}$, $\sigma_{0,\text{upper}}$, $\sigma_{0,\text{lower}}$, $m$, $\mathcal{H}=\{0.9\}$, $\gamma=0.05$, $n_{\text{BO}}=5$, $n_{\text{LBO}}=2$;
@@
11:   $a=a_{3}$$\mathcal{S}_{t}=\{x_{t}^{(k)}\}_{k=1}^{\min(n_{\text{LBO}},i_{\text{max}}+n_{\text{init}}-i)}$;
@@ 4 Experiments
We validated BORA’s performance against current state-of-the-art methods on both synthetic functions and various real-world tasks, with dimensionality ranging from 4 to 15 independent variables. Section 4.1 outlines the experimental setup while Section 4.2 highlights the results. Details on the benchmarks, the method implementations, and the reproducibility details can be found in the SM. The source code is available at https://github.com/Ablatif6c/bora-the-explorer.
@@ 4.1.1 Synthetic Function Benchmarks
  • Levy (10D): A function with a highly rugged landscape. All inputs are bounded in [10,10][-10,10] with the maximum at $1,\dots,1]$.
  • Ackley (15D): A challenging high dimensional function with several local maxima. Input bounds are [30,20][-30,20] with the maximum at $0,\dots,0]$.
@@ 4.1.3 BORA
We implemented BORA using OpenAI’s most cost-effective model at the time, GPT-4o-mini OpenAI (2025), which was not fine-tuned in our effort to make BORA more accessible to users with limited resources. For the BO action implementation, the GP uses a Matérn kernel, and the acquisition function is EI. We set q=5,000q=5{,}000 for $\sigma_{t,\text{mean}}^{\text{GP}}$.
@@ 4.1.4 Baselines
For ColaBO and HypBO, to avoid the impracticality of relying on humans to provide inputs for multiple trials across all experiments, we used the LLM GPT-4o-mini to generate the ‘human’ inputs. Likewise, we used the same task description prompts as used for BORA, to ensure consistency. For HypBO on the Hydrogen Production experiment, we employed the most realistic hypothesis used in Cissé et al. (2024), namely ‘What They Knew’, which encapsulates any human knowledge available prior to the execution of the experiment.
@@ 4.1.5 Experimental Protocol
The optimization performance was measured using the maximum objective value found so far, the cumulative regret, and statistical tests to measure significance. The maximum number of samples was set to 105 to account for realistic budgets with expensive functions. Average results of 10 repeated trials with distinct random seeds are reported. All methods were initialized with $n_{\text{init}}=5$ initial samples apart from LAEA, for which we used 15 initial samples to keep the same number of evaluations to population size ratio as in Hao et al. (2024).
Figure 5: BORA vs Baselines on six experiments. Solid lines show average values while shaded areas indicate $\pm 0.25$ standard error. For visual clarity, some plots are zoomed in to show results up to 60 iterations, as the trends mostly stabilize beyond this point. Full results in SM.
@@ 4.2.1 Synthetic Functions
Figure 3 illustrates the exploration strategy of BORA on the Branin (2D) function. The LLM interventions helped to uncover two out of the three possible locations of the global maximum of Branin. This is further illustrated in Figure 5, which shows that BORA outperforms the baseline comparisons for the higher-dimension functions Levy (10D) and Ackley (15D). A key advantage of BORA is its LLM-informed initial sampling. For mathematical functions, BORA systematically suggests initializing at critical points such as the edges, central points, or other remarkable points like $0,\dots,0]$. This strategy is particularly well-suited for the Levy function, whose search space bounds are symmetric, and its optimum is at $1,\dots,1]$, almost always converging in its initialization stage. However, that strategy is less beneficial on the Ackley function because its bounds are asymmetric. Despite that, the LLM’s ability to reflect and learn from the previous samples appears to help mitigate any unfavorable initializations. While HypBO demonstrates a similar benefit through its initial sampling in hypothesized regions, its performance is comparatively weaker because it relies on random initial sampling within these regions, resulting in less effective exploration of the search space. For ColaBO, which only works with a single input prior, the prior tended to be around one of the edges, which overall limits its convergence speed. Additionally, the left panel of Figure 4 shows how BORA’s iterative hypothesis generation, informed by previous data, helps mitigate stagnant optimization, and discards the LLM when it is no longer needed. Notably, a sharp drop in the GP uncertainty is evident when vanilla BO is used due to is proved exploration-exploitation guarantees, as opposed to the less rigorous LLM where the uncertainty is bound to its inherent sampling strategy. Nevertheless, the dynamic synergistic effect of BO coupled with updated LLM hypotheses allows for faster convergence overall, in comparison to other baselines as illustrated in the last two bar plots of Figure 6.
@@ 4.2.2 Real-World Applications
BORA also exhibits superior performance across diverse real-world optimization problems, following the trends observed in the synthetic benchmarks. As shown in Figure 5, while BORA’s initial sampling is on par with other input-based baselines for the Solar Energy and Sugar Beet Production experiments, its overall performance on all experiments surpasses the baselines significantly. This is particularly evident in the 7D Pétanque experiment, where BORA’s diverse initial hypotheses based on trajectory dynamics led to a remarkable gain in score of 35 in the early stages compared to the baselines. This knowledge and context-based input bridges the knowledge gap typically encountered in early-stage optimization, providing BORA with a critical advantage, as shown in Figure 6. A similar effect is also observed in the Hydrogen Production experiment as illustrated in the right panel of Figure 4. In addition, in the later stage of the optimization the LLM further pushes the optimization to uncover new optima after the progress had stalled, thus gaining more trust. The increasing trend in the GP uncertainty here is a side-effect of the continual interventions of the LLM, which translates to a rather explorative and less exploitative strategy based on its inherent domain reasoning around cumulatively accrued scientific data. This goes beyond some of the near-instant convergence noted in the optimization of the synthetic functions because they are proximal to a special value. Figure 2 illustrates this by showing how the LLM reflects on the progress and generates hypotheses on the Sugar Beet Production experiment. While other baselines, particularly knowledge-based methods such as HypBO and local BO approaches such as TuRBO, demonstrate improved performance as the optimization progresses and more data is gathered, they often struggle to match BORA’s sustained performance as the 105-sample budget mark is approached. In the Hydrogen Production experiment, this adaptive strategy ultimately achieved a 47% reduction in cumulative regret compared to ColaBO, demonstrating BORA’s faster convergence and robustness in navigating complex, high-dimensional search spaces. To assess the significance of the performance difference w.r.t mean cumulative regret between BORA and its best two competitors, we performed a sign test which revealed that BORA performs consistently better than HypBO with a p-value of 0.02 at a 95% confidence level with a Bonferroni correction Bonferroni (1936), but not against ColaBO with a p-value of 0.20, yet still outscoring it in 5 out of 6 tasks. The superior performance of the hybrid approach in BORA was further validated by ablation studies that used only the LLM (action $a_{2}$) for optimizing Hydrogen Production (10D) and the Ackley function (15D) (see SM). While performing quite well in the initial stages for these two problems, the use of the LLM alone was ultimately less effective than the dynamic hybrid BO/LLM approach in BORA. We emphasize that these results do not mean that LLMs are ‘smarter’ than domain experts. Rather, they highlight BORA’s ability to update and refine its hypotheses based on new data, which is not possible in the HypBO implementation Cissé et al. (2024), while also fostering user engagement by generating real-time optimization progress commentary and a final summary report (see SM). One potential limitation of BORA, however, is the stochastic nature of the LLM reasoning, which can diverge considerably even with identical prompts.
@@ A.0.1 Main results
Figure 7: Full results of the BORA vs Baselines on the six experiments. Solid lines show average values while shaded areas indicate $\pm 0.25$ standard error.
@@
To further evaluate BORA’s contributions, we compared it with an LLM-only optimizer, which effectively isolates the LLM component of BORA and functions as an independent optimizer. This comparative analysis allowed us to determine the degree to which the incorporation of Bayesian Optimization (BO) principles enhanced performance beyond the capabilities of the LLM alone. LLM-Only uses identical prompts, with only minor modifications in the role-based prompting mechanism to ensure that it acknowledges its status as the sole optimizer without anticipating BO interventions. It employs the same hypothesis generation mechanism as BORA’s LLM action $a_{2}$, relying exclusively on the outputs of the LLM for optimization. Additionally, LLM-Only uses the same reflection and fallback mechanisms. In the rare event where the LLM fails to produce a valid hypothesis, the corresponding hypothesis point is replaced with a randomly generated point to maintain continuity.
@@
We evaluated BORA and LLM-Only on the more complex synthetic function Ackley (15D) and on the most demanding highest-dimensional real-world task, Hydrogen Production (10D), and reported the maximum objective value found. To ensure comparability with the experimental protocol outlined in the manuscript, LLM-Only was executed with an optimization budget of 105 samples and repeated 10 times.
@@ Results
Figure 8 shows that LLM-Only exhibits competitive performance during the early stages of optimizing both Ackley (15D) and Hydrogen Production (10D). This achievement can be attributed to LLMs’ ability to perform inductive biases, which are used during hypothesis generation. However, LLM-Only fails to sustain its initial performance and quickly stagnates, falling behind BORA significantly in the later stages of the optimization. This stagnation highlights the limitations of relying solely on LLM inference to optimize high-dimensional spaces. In particular, LLMs lack the balanced exploration-exploitation trade-offs that are inherent to BO-based methods, which are achieved through posterior uncertainty modeling. By contrast, BORA, as a hybrid BO/LLM approach, dynamically merges structured sampling from BO with the LLM’s contextual understanding, facilitating sustained improvements as the optimization progresses. Notably, BORA achieved up to a 67% improvement in Hydrogen Production and a 31% improvement in the Ackley benchmark based on the maximum objective values discovered. While BORA considerably outperforms LLM-Only, it is interesting to note that while LLM-Only struggles to sustain its initial, early-stage performance, it still outperforms half of the baselines studied here in this small-budget setting.